I ask som question this morning at about 745am, I received no response. So I'll try again a couple of them I figured out, Still new help. Thanking you in advance. Betty
A. 64 divided 4/2-(5-|2-13|)
B. One less than the product of two numbers is 47...Translate to an equation

C. 56x+14y+282-Combine like terms
D. Simplify 3 divided by 3/2
E. Find the area of opening....
The area of a rectangle with base b and the height is h is bh - A rectangular sports goal is 10ft wide and 8ft high.
Thanking you in advance
Betty

Thanking you in advance
Betty

Answer by stanbon(75887)

A. 64 divided 4/2-(5-|2-13|)
16/2 - (5-11)
= 8 - (-6)
= 8 + 6
= 14
--------------------------
B. One less than the product of two numbers is 47...Translate to an equation
(x*y)-1 = 47
-----------------------
C. 56x+14y+282-Combine like terms
Ans: There are no "like" terms to combine
--------------------------------------------
D. Simplify 3 divided by 3/2
3/(3/2) = 3*(2/3) = (3/3)*2 = 2
-----------------------------
E. Find the area of opening....
The area of a rectangle with base b and the height is h is bh - A rectangular sports goal is 10ft wide and 8ft high.
Ans: area of the opening is 10*8 = 80 sq ft.
-----------
Cheers,
Stan H.
-----------